A man travels 1700 miles on a jet plane and then travels another 300 miles on a propeller plane. The rate of the jet is three times that of the prop. Plane. The total trip took five hrs. What is the rate of the plane?
Answer by dragonwalker(73)

I assume it is asking for the rate of the jet plane and the prop plane.
Let us call the speed of the prop plane x and so as the jet plane is travelling three times as fast its speed with b3 3 times x, otherwise written as 3x
So speed of prop plane = x
Speed of jet plane = 3x
The time it takes for the plane to travel its distance = distance divided by speed.
So:
Time for the jet plane journey = 1700/3x hours
Time for the prop plane journey = 300/x hours
The total time for the journey is 5 hours, so you add the time for each leg of the journey which will equal 5 hours.
This is written as 1700/3x + 300/x = 5
Now solve for x:
Multiply each by 3:
5100/3x + 900/x = 15
Simplify:
1700/x 900/x = 15
Multiply each by x:
1700x/x + 900x/x = 15x
1700 + 900 = 15x (as the x's in each fraction cancel each other out)
2600 = 15x
divide each side by 15 to find x:
2600/15 = 15x/15
173 1/3 = x
so the speed of the the prop plane is 173 1/3 miles per hour
and multiply this by 3 (as the speed of the jet plane is 3 times the speed of the prop plane) to find the speed of the jet plane = 520 miles per hour